Collins 51J-4 sn 3516 

 


The cabinet arrived 3/23/06

 

It was packed in a double walled microwave oven box, with some of the microwave foam edges, not an exact fit, but worked OK.  The cabinet was able to shift inside the box, but no damage noted.  I think the whitish marks on the top of the cabinet are a result of movement, and may come clean with some soap/water.

 

The camera & lighting make color rendition a bit flakey.  It is a reasonable "St. James Gray".

It has been over-painted, probably recently.  Areas of scratches thru the original paint, and even a few fairly large areas, can be seen in the close-ups, but from a little distance it doesn't look bad.


Receiver arrived 4/26/06

packed by UPS Store

51J4pix/pkg03F.JPG (43651 bytes)   51J4pix/pkg08F.JPG (49547 bytes)  51J4pix/pkg10F.JPG (59963 bytes)  51J4pix/pkg12F.JPG (54025 bytes)  51J4pix/pkg16F.JPG (66061 bytes)

   Top of box was depressed about an inch, probably by something placed on top of it.  The UPS driver and I opened the top to inspect, saw that it was double boxed, with wadded paper padding, looked OK to us.  It was well packed, no damage found when totally unpacked.


51J4pix/arrd01.JPG (68473 bytes) 51J4pix/arrd03.JPG (66758 bytes) 51J4pix/arrd05.JPG (78494 bytes) 51J4pix/arrd06.JPG (78927 bytes) 51J4pix/arrd07.JPG (73150 bytes) 51J4pix/arrd08.JPG (49322 bytes) 51J4pix/arrd09.JPG (68102 bytes) 51J4pix/arrd10.JPG (84043 bytes) 51J4pix/arrd11.JPG (73959 bytes) 51J4pix/arrd12.JPG (82004 bytes) 51J4pix/filt1.JPG (62261 bytes) f51J4pix/ilt2.JPG (59940 bytes)

"as received" observations

   Dirty inside & panel.  Has bottom cover (missing 7 screws), no top cover or tools.  Has all knobs, no cracks or chips noted, but they may have been clear-coated, which is flaking off.  That may polish out.  Has 7 IERC type tube shields, 5 standard tube shields, and is missing 6.  PTO is a 70E-15, s/n 5712, the seal may not have been broken.  The band indicator drum is good, not badly yellowed, 1 or 2 small scratches not thru the paper, 1 scratch at rt end thru the paper, made by the lamp holder, not visible from the front.  Has both hold-down clamps, for the rectifier/filter can and the crystal.  Under chassis is fairly clean, no obvious previous work has been done.  There may have been a glued ID tag on the panel to the left of the Collins s/n tag, the paint has been touched up there, and is not wrinkle.  The 2-section filter cap is dated 6-59, ESR checks 0.6-0.8 for both secns, fair to good, but will re-stuff the can.  Has "vintage" 2-wire line cord, secured with hot-melt glue at the cable clamp.  Will replace with proper 3-wire cord.  Only one IF filter (3 kc) is installed, and extra (1.400kc) has been supplied for possible use.

   Powered up with variac fairly quickly as it had been under power recently, draws less than 1 amp.  Dial lights are all dim at full AC input.  PTO is off 8-9 kc end-to-end, will be remedied.  S-meter zero was at about 10db at first, after about 5 minutes dropped below zero to the peg, and the output volume dropped.  It did not change back the next day at power up.  May be the "notorious C-204".  The sensitivity drops on the higher bands, but they all are operable.  24-5 - 25.5 and 25.5 - 26.5 bands are approx. 15kc off compared to the others, will ck to see if they use the same band xtal. 


Tubes were removed and initial cleaning performed.  It's still pretty cruddy in the area around the power transformer.

ptclean2.JPG (68640 bytes)    ptclean1.JPG (80397 bytes) 

Some electrical inspection and checking performed, ESR checked for the 4 bathtubs that contained electrolytic capacitors.  2 were OK, one marginal and one bad.  These, and the others containing paper caps, will be restuffed.

All knobs were removed to allow the front panel to be removed & left hanging by it's cables, to allow more cleaning and PTO removal.

ptclean3.JPG (74162 bytes) 

 The PTO was removed, one turn removed from the compensating coil to improve the end to end span, which was about 8kc off. 

PTO3.JPG (79266 bytes)   PTO4.JPG (78729 bytes)

Pix show the original (left) & replacement (rt) feedback cap, and the dessicant capsules b4 & after removal.

~1 turn was removed from the compensating coil.  After about 8 iterations of taking readings every 100kc point and adjusting the compensating slug,  excellent results were had.  From  end to end, 2000  thru 3000kc, no point was more than 1/2 kc off, most within 1/4.

dial turns   -1/2            0             1                2                3               4            5                6                7            8                9               10           10-1/2

freq.       1950.57    2000.00    2100.03    2200.30    2400.20    2500.27    2600.16    2700.19    2800.03    2899.59    2999.66    3000.01    3047.00

This was achieved after "soaking" the PTO, leaving the filaments on for several days.  It may be slightly different with less "warmup".

I've given the PTO cover can a light buffing, ready to reinstall it and the PTO.


Cleaning this week

Dialdirt.JPG (62419 bytes) dirt on rag from one side of the KC dial

arrd04r.jpg (219295 bytes) FrtPnlB401.JPG (67495 bytes) FrtPnlB402.JPG (76379 bytes) FrtPnlB403.JPG (83933 bytes) panel before cleaning

 FrtPnlmid1.JPG (63408 bytes) FrtPnlmid2.JPG (65297 bytes) panel after 2 passes, need to clean up a bit still.  The wrinkle had been partially filled with dirt, nicotine, etc.  Note, no red "7" any more.

Yuk01.JPG (69854 bytes) Yuk02.JPG (82910 bytes) Yuk03.JPG (86516 bytes)  Yuk05.JPG (71636 bytes) Yuk07.JPG (67871 bytes) Yuk04.JPG (60274 bytes)

grunge near power transformer, very sticky, not sure if it's tar from the transformer (don't think so, and hope not), or something that was spilled inside that end.  Decided to completely remove the side panel in order to do a better cleaning job, only about 2 dozen screws holding it & parts mounted to it.  May not be able to do too much down between the filter unit & xfmr.

cansB402.JPG (48794 bytes)  I took the IF cans off & found the whole problem ;-)    - the engineer was dead.    DeadEngr.JPG (32551 bytes)

NoYuk01.JPG (76814 bytes) NoYuk02.JPG (82972 bytes) NoYuk03.JPG (81404 bytes) NoYuk04.JPG (82868 bytes) NoYuk05.JPG (89295 bytes)

After cleaning the grunge.  Last pic is the bottom of the power transformer, I don't think the stuff came out of it.

Bktogether1.JPG (81141 bytes) Bktogether2.JPG (74194 bytes) Bktogether3.JPG (79303 bytes)

IF cans back in place, after buffing.  Starting to look pretty.


lampline.JPG (66470 bytes)  I had noted tape over wiring at the meter earlier.  Apparently the 6.3vac line to the dial lamps had been shorted to ground at one of the lamps (I've heard of that problem before), and the wire had burned it's insulation off.  A previous owner had spliced a new wire on at the point near the meter where the wire comes out of a cable, and covered it with tape.  I took it loose and put heat shrink tubing over the splice.  

PDwiring01.JPG (77780 bytes) PDwiring02.JPG (78543 bytes)  Before re-installing the panel I re-installed the PTO and installed the cables necessary for switching in a product detector, which will be added.  The black wires are small dia. coax. cable for the new lines.  All knobs were cleaned and waxed, as was the front panel dial bezel.  I had noted earlier that all the dial lamps were dim, that's because they are 14.4v lamps and the supply is 6.3vac.

   All tubes were tested, one 6BE6 was very weak and was replaced.  All tube pins were straightened and DeOxit applied as they were re-installed.  I replaced the 3 9-pin tube shields that were missing with IERC heat dissipating shields like the ones that had been installed on many of the 7-pin tubes.  Several 7-pin shields are still needed, I do not have any more IERC's of that size.

ACcord.JPG (89320 bytes)  A 3-wire power cord was installed, along with proper 250vac rated X-Y bypass capacitors.

   A 10 watt wire wound resistor can be seen standing up from the panel just to the right of the new orange caps.  It's an addition by a previous owner, and is faulty, open near one end where a tap is for a bias/AVC line.  The bias voltage was measured when the rcvr was powered up after re-installing the front panel and all tubes, and was found to be -70 vdc.  It should be between -60 and -56.  This is a common problem on the R-388/51J-3,4 receivers, partly due to the increased AC line voltage we now have.  There are a couple of minor changes that can be made to bring it more into line, plus I will have to trace out the wire that is on the extra resistor tap, and remove or replace the resistor.  The low sensitivity is at least partially due to the high negative voltage on the AVC.  The calibrator signal is strong on the lower bands.  The sensitivity changes abruptly after a few minutes of warmup, and does change back and forth some for another few minutes before just staying poor.

   There is an intermittent "crackling" or noise, which I can't make to be related to anything, might be a tough one to chase down.  Before I removed the bathtub caps, shown below, I did replace C-204 & C-226, 100pf micas, in the AVC section, hoping that might cure the intermittent stuff, it didn't.  I've seen similar symptoms which were due to a bad C-204.


C-204 is a mica coupling capacitor at the AVC detector, it is a common failure item.

V110C204r.jpg (209082 bytes)  V110C204epl.JPG (94455 bytes) C-204 & C-226 original & replaced.  The black loop is a cable for the prod. det. installation, not connected at this end yet.

Btubsoutr.jpg (270721 bytes)  5 bathtubs removed and cleaned out    5stuffed.JPG (74499 bytes)  5 of 7 re-stuffed

7backinr.jpg (315280 bytes)  All 7 back in   There are 7 bathtubs, with a total of 11 sections; 4 electrolytics, 7 bypasses, paper originally, now polyester.

   I was hoping that replacement of the bathtubs would result in 2 things, better bias voltage, and no more crackling/change in sensitivity.  The bias voltage is developed by voltage drop over resistors in the center tap of the B+ winding of the power transformer.  That is dependent upon the B+ current draw, which hopefully would come down some with new bypass capacitors.  It did slightly, from -70 to -68. Still not low enough, and that will be fixed.  It still has some crackling and change in sensitivity, but in preliminary alignment adjustments it may be that some of that is caused by loose or broken wiring at L-115, in the first IF.  That may be the one that had the slug which had fallen out.  Overall sensitivity is still a bit low on band 1, low BC band, and probably the others.  I may have to resort to signal tracing with the scope to chase these down.


Most of the "crackling sound" was found to be caused by a 6BA6 tube in the filter module, V-302.  It had tested good, but had not been tested for noise.  Some of the changes in sensitivity, and some crackling was also due to a bad internal ground connection to the external ground post in the filter module.  A previous owner had added an external jumper wire from the top of the module to the top of the xtal filter can, as seen in one of the early pix.  I installed a hidden ground strap below, to avoid the difficult job of removing the filter module.  There is no secure mechanical connection between the filter module and the chassis, as originally designed.  The mechanical filter module was originally designed by Collins as an "kit" addition to the R-388/51J-3 model, to add mechanical filters to a receiver that did not have them originally.  When they produced the 51J-4, to include the mechanical filters, the merely used the additional box, as if it was an afterthought modification, and did not re-configure the chassis.

A product detector is to be installed, basically the circuit of Dallas Lankford, similar in many ways to those of Orr & Lee.  The 6BA6 BFO tube, V-114, is replaced by a 6BE6 which is used as the BFO oscillator and a product detector.  The original diode detector is retained, so AM reception is not affected.  

PDstrip1.JPG (91502 bytes) PDstrip2.JPG (93360 bytes) Components stripped out for product detector installation, one wire from T-104 has been relocated on V-110.

PDin2.JPG (92106 bytes) Product detector almost complete, at V-114.

It was necessary, as noted by Lankford and others, to custom select values for the coupling capacitor and shunt resistor that feed the IF signal to the product detector.  This is made necessary due to differences in individual receivers.  He had found that a 3mmf coupling cap and 8200 ohm resistor were optimum for him.   5mmf & 15k were necessary here, not inconsistent with what others have done.  This is done to match the audio output levels when using the prod. det. vs. the AM diode det.  Audio level was insufficient using Lankford's original values.

Some modifications to the AGC are also needed, the attack time is reduced to eliminate "popping" on CW reception, and the hang time is increased to eliminate, or reduce, "pumping".

PDin4.JPG (80831 bytes)  some of the AGC mod. at the AGC amp. tube, V-110, next to V-114, the prod. det./BFO, before addition of the 1.5mf "hang time" cap.  Lee's attack time mod. was used, with a 820k ohm resistor replacing R144, and bridged with a 2.2k res. & diode.

PD_AGCcompl2.JPG (87819 bytes) PD_AGCcompl1.JPG (90831 bytes)  Close-up and overall shots of the completed AGC & Prod. Det. installation.   It pretty much covers up V-114, AGC amp., and V-110, prod. det/BFO tube sockets. 

Here's the Prod. Det. & AGC info   51J4sn1.gif (1167163 bytes)

With things all back together and operating well, I still needed to chase down the low sensitivity problem.  Compared to the "good" 51J-4, broadcast band signals, for example, were some 20-30 db less on this receiver.  I replaced the 1st RF tube, a 6AK5, with a new one, and it helped on higher bands, but only slightly.  Following the schematic, and checking components, I finally suspected the standby relay, which switches the antenna from the 1st RF stage to ground when on stdby.  After cleaning the contacts with a strip of paper soaked with DeOxit, the receiver came alive.  This relay is never operated unless it is connected externally.  Occasional operation would keep the contacts clean.

I  have put the new aftermarket 6kc filter in, had to put solder on the pins.  The provider didn't do a real nice job for pins - he used small dia. threaded brass screws or rod, abt #2 or #3.  Just too small.  He had already put a little solder right at the tips, just abt enuf to fill in the threads.  Needed a little more, in dia. and further down. It does make BC signals sound nicer than the 3.1kc stock filter.


I started to do some touch up on the front panel scratches, but my small bottle of paint had dried up and is gooey.   The next day I used some of my St. James Gray cabinet paint, much nicer.

Have put the bottom cover back on, with all screws, will do a final alignment with it in place. 

Then I think this one's done.

------------------------

The above have been done, touchup looks much better.                                                      and as a reminder of earlier stuff,

FrtPnlB4LL01.JPG (47547 bytes)  PnlLLtur.jpg (238424 bytes)           FrtPnlB4Tag01.JPG (50416 bytes)  PnlTLtur.jpg (278144 bytes)                 Yuk03.JPG (86516 bytes) NoYuk01.JPG (76814 bytes)

        before                    after                             before                   after                                before                   after


Here's the minor mod. needed at the back of the cabinet to let the rcvr fit in.  Notch out skirt to clear the studs for the top cover.  I think the cabinet is for a 75A3 receiver.

notched.jpg (173020 bytes)